OpenStudy (anonymous):

Multiply x to the 2 fifths power times x to the 2 ninths power

OpenStudy (anonymous):

pls help

Elsa213 (elsa213):

To multiple powers you add them 2/5+2/9 = 18/45+10/45=28/45 x^(18/45) https://answers.yahoo.com/question/index?qid=20150209154454AAwfqUk

Elsa213 (elsa213):

c:

OpenStudy (anonymous):

thank you so much !!

OpenStudy (anonymous):

\[x^\frac{ 2 }{ 5}\ \times\ x^\frac{ 2 }{ 9}= x^(\frac{ 2}{5}+\frac{ 2 }{ 9 })\\=x^\frac{ 28 }{ 45}\]

OpenStudy (anonymous):

@Elsa213 x^(28/45) not x^(18/45). common mistake in arithmetic operations.

Elsa213 (elsa213):

I'm sorry, but I was showing him an example of what to do. c:

OpenStudy (anonymous):

to err is human! sign of intelligent ones.

Elsa213 (elsa213):

;-;